OpenStudy (anonymous):

A separable differential equation is a first-order differential equation that can be algebraically manipulated to look like: A) f(y)dy = g(x)dx B) f(x)dx + f(y)dy C) g(y)dx = f(x)dx D) f(x)dx = f(y)dy E) Both f(y)dy = g(x)dx and f(x)dx = f(y)dy

myininaya (myininaya):

separation or variables mean to get your x's dx on one side and your y's dy on the other

myininaya (myininaya):

which one says that

OpenStudy (anonymous):

E

OpenStudy (anonymous):

E?

myininaya (myininaya):

Yes

myininaya (myininaya):

E=A+D :)

OpenStudy (anonymous):

It was actually A. thanks anyway. :)

OpenStudy (anonymous):

yes E because in both case, we have a differential equation in form f(y)dy/dx=g(x) a f(x)dx/dy=g(y) and both can be algebraically manipulated to become f(y)dy=g(x)dx and f(x)dx=g(y)dy

myininaya (myininaya):

You can have f(*)=*^2 f(x)=x^2 f(y)=y^2 x^2 dx=y^2 dy this is still done by separation or variables it is totally E

OpenStudy (anonymous):

I submitted the answer and the computer program told me A. you want to see there explanation? here you go:

OpenStudy (beginnersmind):

Yeah, either the program is wrong or you copied the options incorrectly. E is definitely right.

OpenStudy (anonymous):

I guess so.....

OpenStudy (beginnersmind):

although f(x)dx=f(y)dy is a very specific equation

myininaya (myininaya):

If it said all first order equations that are separable can be written in the form A) f(y)dy = g(x)dx B) f(x)dx + f(y)dy C) g(y)dx = f(x)dx D) f(x)dx = f(y)dy E) Both f(y)dy = g(x)dx and f(x)dx = f(y)dy I would have said A But it said "a"